I need help answering the following math problem: Going into the final exam, which will count as two-thirds of the final grade, Mike has test scores of 86, 80, 84, and 90. What score does Mike need on the final in order to earn a B, which requires an average score of 80? What does he need to earn an A, which requires an average of 90?
I began answering the first part of the problem by using the following equation and solving for x: 86 + 80 + 84 + 90 + 2/3x = 340 + 2/3x. X = 90. So, Mike needs a score of 90 to earn a B and using the same formula, he needs a score of 165 to earn an A.
However, I am not sure if I solved this problem correctly. Please advise where I might have messed up. Thank you in advance.

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW.
Going into the final exam, which will count as two-thirds of the final grade
FIRSTLY WHAT IS MEANT BY THIS STATEMENT?IT IS NOT CLEAR..DOES IT MEAN 2/3 RD.OF MARKS GOT IN THE FINAL EXAM WILL BE COUNTED?IT DOES NOT MAKE SENSE...IF IT HAS NO RELEVANCE TO THE PROBLEM IT IS OK......
ORELSE THE ONLY INTERPRETATION I CAN THINK OF IS THAT THE LAST EXAM IS FOR A TOTAL OF MAXIMUM 150 MARKS AND HENCE TO BRING IT DOWN TO A MAXIMUM OF 100 MARKS AVERAGE AS IS THE CASE WITH OTHER EXAM MARKS , WE NEED TO MULTIPLY THE MARKS GOT IN THE FINAL EXAM BY 2/3..
Mike has test scores of 86, 80, 84, and 90. What score does Mike need on the final in order to earn a B, which requires an average score of 80?
OK WITH THIS THE PROBLEM IS COMPLETE AND CAN BE SOLVED AS SHOWN BELOW..
What does he need to earn an A, which requires an average of 90?
THIS IS A SECOND PART OF THE PROBLEM WHICH CAN BE SOLVED IN SAME WAY..
I began answering the first part of the problem by using the following equation and solving for x: 86 + 80 + 84 + 90 + 2/3x = 340 + 2/3x.
THIS IS NOT THE EQN..THE EQN. IS
340+2X/3= 5*80=400...OR 340+X=400 ,IF WE NEGLECT THAT 2/3 PART OF THE QUESTION AS I SAID ABOVE.
SEE THE PROBLEM SOLVED BY ME EARLIER (GIVEN BELOW)TO COMPARE YOUR WORKING WITH MINE.OF COURSE AS I DEDUCED LATER YOUR WORKING MUST BE APPARENTLY SAME AND OK... .THE ANSWER IS X=60..AND IF THAT 2/3 FACTOR IS REALLY WHAT YOU MEANT THEN YOUR ANSWER OF X=90 IS CORRECT SINCE IF ONLY 2/3 OF THE SCORE IS TAKEN WE GET 90*2/3=60
X = 90.
So, Mike needs a score of 90 to earn a B
CORRECT!!! IF THE ABOVE INTERPRETATION IS CORRECT..
and using the same formula, he needs a score of 165 to earn an A.
CORRECT!!! IF THE ABOVE INTERPRETATION IS CORRECT..
However, I am not sure if I solved this problem correctly.
HAVE CONFIDENCE IN YOUR ABILITY!YOU HAVE DONE EXTREMELY WELL AND BOUND TO DO BETTER WITH CONFIDENCE..
Please advise where I might have messed up.
NO NO NO....A THOUSAND TIMES..YOU HAVE DONE EXCEPTIONALLY WELL..DONT FEEL THAT WAY AT ALL ...YOU ARE ON YOUR WAY TO BECOME AN EXPERT!!!
Thank you in advance.


SEE THE FOLLOWING EXAMPLE...
Judy needs a mean(average) socre of 86 on four tests to earn a midterm grade of B. If the the mean of her first three test was 83, what is the lowest score ona 100-point scale that she can receive on the fourth test to have a midterm grade of B? **Let 't' represent the unknown test grade
1 solutions
--------------------------------------------------------------------------------
Answer 14973 by venugopalramana(912) on 2006-02-15 23:37:26 
Judy needs a mean(average) socre of 86 on four tests to earn a midterm grade of B. If the the mean of her first three test was 83, what is the lowest score ona 100-point scale that she can receive on the fourth test to have a midterm grade of B? **Let 't' represent the unknown test grade
TOTAL TESTS =4
AVERAGE NEEDED TO GET B GRADE = 86
SO TOTAL MARKS NEEDED IN 4 TESTS =AVERAGE *NUMBER OF TESTS=86*4=344
NUMBER OF TESTS ATTENDED BY JUDY =3
AVERSGE GOT IN 3 TESTS =83
SO TOTAL MARKS GOT IN 3 TESTS =AVERAGE * NUMBER OF TESTS=83*3= 249
SO MARKS TO BE GOT IN 4 TH. TEST =344-249= 95